# Balancing properties of tropical moduli maps

### v1

@@ 3. Main theorem
By definition, the underlying graph of $\mathrm{trop}({\mathcal{C}}_{\eta})$ is the dual graph $\mathbb{G}_{s}$ of $\widetilde{\mathcal{C}}_{s}$. We first show that $\mathbb{G}_{s}$ only depends on the stratum $S$. Let $z$ be a node of $\widetilde{\mathcal{C}}_{s}$. Since ${\mathcal{C}}^{0}$ is pulled back from $\overline{M}$, there exist an étale neighborhood $V$ of $s$ in $B^{0}$ and a function $g_{z}\in\mathcal{O}_{V}(V)$ vanishing at $s$ such that the family ${\mathcal{C}}^{0}\times_{B^{0}}V$ is given by $xy=g_{z}$tale locally near $z$. After shrinking $V$, we may assume that the latter is true in a neighborhood of any node of $\widetilde{\mathcal{C}}_{s}$, and that the morphism $V\to B^{0}$ factors through $U\subseteq B^{0}$. By assumption, ${\mathcal{C}}$ is smooth over $B\backslash H$, hence each $g_{z}$ is invertible on the complement of $D_{B^{0}}$. However, $V$ is normal, and $g_{z}(s)=0$ for all nodes $z$, hence each $g_{z}$ must vanish along some component of $D_{B^{0}}\cap V$ that intersects $S$, and therefore contains $V\cap S$. Thus, all the $g_{z}$’s vanish identically along $S$. This implies that the dual graph $\mathbb{G}_{s}$ is étale locally constant along $S$. Moreover, since $\widetilde{\mathcal{C}}_{S}\rightarrow S$ is split, the identification of $\mathbb{G}_{s}$ with the dual graph of the fiber over the generic point of $S$ is canonical. Therefore, $\mathbb{G}_{s}$ only depends on $S$. Set ${\mathbb{G}}_{S}:={\mathbb{G}}_{s}$. Then for any $S\leq T\in\mathrm{Str}(B^{0},H^{0})$ there is a natural contraction $\phi_{S,T}\colon{\mathbb{G}}_{S}\rightarrow{\mathbb{G}}_{T}$ obtained by the specialization of the curve over the generic point of $T$ to the curve over the generic point of $S$.
@@
where $k_{i}:=n_{i}+\frac{1}{l}n_{\pi}=\frac{1}{l}\mathrm{ord}_{D_{i}}(g_{z})$ if $i\leq a$ and $k_{i}:=n_{i}=\mathrm{ord}_{D_{i}}(g_{z})$ otherwise. We claim that the $k_{i}$’s are independent of the choice of $s\in S$. Indeed, for another $s^{\prime}\in S$, the corresponding $g_{z}$ and $g^{\prime}_{z}$ differ by an invertible function on the intersection of the corresponding étale neighborhoods of $s$ and $s^{\prime}$, and hence also at the generic points of the $D_{i}$’s for $0\leq i\leq a+b$. Thus, $\mathrm{ord}_{D_{i}}(g_{z})=\mathrm{ord}_{D_{i}}(g^{\prime}_{z})$ and therefore $k_{i}=k^{\prime}_{i}$ as claimed. Furthermore, we see that the length function $\ell(\gamma)$ depends only on the tropicalization $q:=\mathrm{trop}(\eta)=\left(\nu\left(\varphi^{*}(x_{0})(\eta)\right),\dotsc,\nu\left(\varphi^{*}(x_{a+b})(\eta)\right)\right)$ rather than the point $\eta$ itself.
@@
$-\mathrm{ord}_{\widetilde{\mathcal{C}}_{s,u}}\left(f^{*}(x^{m})|_{{\mathcal{C}}_{\overline{\eta}}}\right)=-n_{\pi}-\sum_{i=0}^{a+b}n_{i}\cdot\nu\left(\varphi^{*}(x_{i})(\eta)\right)=-n_{\pi}-\sum_{i=0}^{a+b}n_{i}y_{i},$
@@
In the first two steps, we have constructed the data of a family of parameterized topical curve over $\Lambda$ the extended degree, the combinatorial type for each face of $\Lambda$, the weighted contractions, and the integral affine functions $\ell_{S}$ and $h_{S}$. Furthermore, we have verified the compatibility conditions of Definition 2.7, and have seen that the resulting family of parameterized topical curves over $\Lambda$ satisfies assertion (1).
Proof of Part (2) of Theorem 3.1
@@
Let $\{S_{i}\}$, where $a+b+1\leq i\leq a+b+r+t$, be the set of strata of codimension one in $\overline{S}$. Each $S_{i}$ corresponds to a face $W_{i}\in\mathcal{F}(\Lambda)$ containing $W$ as a codimension-one face. Suppose $\overline{S}_{i}=\overline{S}\cap D_{i}$ for some component $D_{i}$ of $D_{B^{0}}$. We may further assume that $D_{i}$ is vertical for $a+b+1\leq i\leq a+b+r$, and horizontal for $a+b+r+1\leq i\leq a+b+r+t$. Note that the $D_{i}$’s are uniquely determined by the $S_{i}$’s but may not necessarily be distinct. In the sequel, we will mainly consider the following components of $D_{B^{0}}$, which have non-empty intersection with $\overline{S}$: the components $D_{0},\dotsc,D_{a+b}$, which are distinct and contain $\overline{S}$; and the components $D_{a+b+1},\dotsc,D_{a+b+r+t}$ constructed above.
@@
Since $\overline{S}$ is étale locally isomorphic to an affine space and the $\overline{S}_{i}$’s correspond to the coordinate hyperplanes, we can pick a (possibly reducible) curve $C$ in $\overline{S}$ that has positive intersection numbers with each $\overline{S}_{i}$ and none of whose components is contained in $\cup\overline{S}_{i}$. In particular, when $\dim S=1$ (equivalently, $W$ has codimension one in $\Lambda$), we set $C:=\overline{S}$. Set $l_{i}:=l$ for $i\leq a$, and $l_{i}:=1$ for $a+1\leq i\leq a+b$ and for $i>a+b+r$. Then (3.2) would follow from the following equality
@@
$-\sum_{i=0}^{a+b}l_{i}k_{i}D_{i}\cdot C+\sum_{i=0}^{a}l_{0}k_{0}D_{i}\cdot C=-\sum_{i=0}^{a+b}l_{i}k_{i}D_{i}\cdot C+\sum_{i=0}^{a}lk_{0}D_{i}\cdot C.$

### v2

@@ 3. Main theorem
[…]
@@
where $k_{i}:=n_{i}+\frac{1}{l}n_{\pi}=\frac{1}{l}\mathrm{ord}_{D_{i}}(g_{z})$ if $i\leq a$ and $k_{i}:=n_{i}=\mathrm{ord}_{D_{i}}(g_{z})$ otherwise. We claim that the $k_{i}$’s are independent of the choice of $s\in S$. Indeed, for another $s^{\prime}\in S$, the corresponding $g_{z}$ and $g^{\prime}_{z}$ differ by an invertible function on the intersection of the corresponding étale neighborhoods of $s$ and $s^{\prime}$, and therefore also at the generic points of the $D_{i}$’s for all $0\leq i\leq a+b$. Thus, $k_{i}=\mathrm{ord}_{D_{i}}(g_{z})=\mathrm{ord}_{D_{i}}(g^{\prime}_{z})=k^{\prime}_{i}$ as claimed. Furthermore, we see that the length function $\ell(\gamma)$ depends only on the tropicalization $q:=\mathrm{trop}(\eta)=\left(\nu\left(\varphi^{*}(x_{0})(\eta)\right),\dotsc,\nu\left(\varphi^{*}(x_{a+b})(\eta)\right)\right)$ rather than the point $\eta$ itself.
@@
$\mathrm{ord}_{\widetilde{\mathcal{C}}_{s,u}}\left(f^{*}(x^{m})|_{{\mathcal{C}}_{\overline{\eta}}}\right)=-n_{\pi}-\sum_{i=0}^{a+b}n_{i}\cdot\nu\left(\varphi^{*}(x_{i})(\eta)\right)=-n_{\pi}-\sum_{i=0}^{a+b}n_{i}y_{i},$
@@
In the first two steps, we have constructed the data of a family of parameterized tropical curve over $\Lambda$ - the extended degree, the combinatorial type for each face of $\Lambda$, the weighted contractions, and the integral affine functions $\ell_{S}$ and $h_{S}$. Furthermore, we have verified the compatibility conditions of Definition 2.7, and have seen that the resulting family of parameterized tropical curves over $\Lambda$ satisfies assertion (1).
3.2. Proof of Part (2) of Theorem 3.1
@@
Let $\{S_{i}\}_{i=a+b+1}^{a+b+r+t}$ be the set of strata of codimension one in $\overline{S}$. Each $S_{i}$ corresponds to a face $W_{i}\in\mathcal{F}(\Lambda)$ containing $W$ as a codimension-one face. Suppose $\overline{S}_{i}=\overline{S}\cap D_{i}$ for some component $D_{i}$ of $D_{B^{0}}$. We may further assume that $D_{i}$ is vertical for $a+b+1\leq i\leq a+b+r$, and horizontal for $a+b+r+1\leq i\leq a+b+r+t$. Note that the $D_{i}$’s are uniquely determined by the $S_{i}$’s but may not necessarily be distinct. In the sequel, we will mainly consider the following components of $D_{B^{0}}$, which have non-empty intersection with $\overline{S}$: the components $D_{0},\dotsc,D_{a+b}$, which are distinct and contain $\overline{S}$; and the components $D_{a+b+1},\dotsc,D_{a+b+r+t}$ constructed above.
@@
Since $\overline{S}$ is étale locally isomorphic to an affine space and the $\overline{S}_{i}$’s correspond to the coordinate hyperplanes, we can pick a (possibly reducible) curve $C$ in $\overline{S}$ that has positive intersection numbers with each $\overline{S}_{i}$ and none of whose components is contained in $\bigcup\overline{S}_{i}$. In particular, when $\dim S=1$ (equivalently, $W$ has codimension one in $\Lambda$), we set $C:=\overline{S}$. Set $l_{i}:=l$ for $i\leq a$, and $l_{i}:=1$ for $a+1\leq i\leq a+b$ and for $i>a+b+r$. Then (3.2) would follow from the following equality
@@
$\sum_{i=0}^{a+b}l_{i}k_{i}D_{i}\cdot C+\sum_{i=0}^{a}l_{0}k_{0}D_{i}\cdot C=-\sum_{i=0}^{a+b}l_{i}k_{i}D_{i}\cdot C+\sum_{i=0}^{a}lk_{0}D_{i}\cdot C.$